OpenStudy (anonymous):

Look at the following numbers: −3, −1, 0, 3 Which pair of numbers has a sum of 0? 0, −1 −3, 3 −1, 3 −1, 0

OpenStudy (welshfella):

which 2 numbers only differ by the sign?

OpenStudy (anonymous):

-3 -1

OpenStudy (welshfella):

no - these 2 have the same sign and number part is different

OpenStudy (anonymous):

O_o

OpenStudy (anonymous):

dont get it

OpenStudy (welshfella):

recall that a number without a sign is positive 2 is the same as +2

OpenStudy (anonymous):

okay

rvc (rvc):

Suppose x is the number then other number such that their sum equals 0 will be -x x-x=0

OpenStudy (anonymous):

-1,0

OpenStudy (anonymous):

is that it

OpenStudy (welshfella):

no because 0-1 = -1 you need something like 2 - 2 = 0

OpenStudy (anonymous):

im so stupid -3,3

OpenStudy (welshfella):

right! 3 - 3 = 0

rvc (rvc):

right answer

OpenStudy (anonymous):

thank you guys for your help

OpenStudy (welshfella):

yw

rvc (rvc):

All the best.